# Unifying the Perspectives of NLP and Software Engineering: A Survey on Language Models for Code

### v3

@@ 5.6 Diffusion Models
Currently the Transformer architecture dominate text generation, but several works (Li et al., 2022c; Lin et al., 2023) have also adopted Diffusion Models (Ho et al., 2020) from computer vision for text generation. Recently CodeFusion (Singh et al., 2023) also introduces diffusion models into code modeling, and demonstrates that a 75M diffusion model can outperform StarCoder, CodeT5+, and GPT-3 on three code synthesis datasets.
@@ 5.7 Instruction Finetuning and Reinforcement Learning for Code
In natural language processing, training models on a diverse set of tasks with instruction prefix, known as instruction finetuning, has been shown to unlock the ability of cross-task generalization (Ouyang et al., 2022; Chung et al., 2022; Iyer et al., 2022). At first, these instruction data samples are manually compiled or crowd-sourced (Wei et al., 2022b; Sanh et al., 2022), but later researches find LLM-generated instructions to be sufficient (Wang et al., 2023c; Honovich et al., 2023).
Following these works in natural language, researchers from the code community have applied instruction tuning to their models as well. Wang et al. (2023d) finetune CodeT5+ with 20K instruction data generated by InstructGPT (Ouyang et al., 2022) to obtain InstructCodeT5+. WizardCoder (Luo et al., 2023) follows the methods of WizardLM (Xu et al., 2023a) to evolve 20K code Alpaca (Taori et al., 2023) samples into a 78K dataset and uses it to finetune StarCoder. Pangu-Coder 2 (Shen et al., 2023) also uses WizardLM’s Evol-Instruct to generate 68K instruction samples from 20K code Alpaca, but also introduces reinforcement learning via Rank Responses to align Test & Teacher Feedback (RRTF). OctoCoder (Muennighoff et al., 2023), on the other hand, takes a different path and uses Git commit histories as instruction data to finetune StarCoder and CodeGeeX2. More recently, CodeFuse (Di et al., 2023) also employs multitask-finetuning and explicitly introduces multiple downstream tasks into their instruction data. The performance of these instruction finetuned code models can also be found in Table 4.
@@
CodeRL (Le et al., 2022) is one such model, which defines four levels of rewards for each generated program (viz. compile error, runtime error, unit test failure, pass) as well as fine-grained token-level reward estimated by a critic model. The actor model, which is an extention of CodeT5, is then trained with REINFORCE algorithm (Williams, 1992). Similarly, CompCoder (Wang et al., 2022c) and PPOCoder (Shojaee et al., 2023) train CodeGPT and CodeT5 respectively with proximal policy optimization (Schulman et al., 2017), while RLTF (Liu et al., 2023d) proposes fine-grained feedback based on the error information and location provided by the compiler, as well as adaptive feedback that takes the ratio of passed test cases into account.
@@ 7.1 LLMs Extended with Coding Tools
Researches in the NLP community have shown that LLMs can learn to use external tools such as calculators, MT systems, and search engines (Thoppilan et al., 2022; Schick et al., 2023). As such, interpreter has been used to augment LLMs in complex reasoning tasks. PAL (Gao et al., 2023b) and PoT (Chen et al., 2022) both extend Codex with Python interpreters for numerical calculations, while ViperGPT (Surís et al., 2023) extends it further by calling vision APIs to extract information from visual input and answer related questions.
Apart from alleviating the burden of numerical calculation in abstract reasoning tasks, interpreter also provides feedback on the process of code generation itself, together with unit tests. CodeT (Arteca et al., 2022) and TiCoder (Bareiß et al., 2022) use Codex to generate unit tests, which are run against generated code samples to improve the model’s performance on code synthesis. Similarly, TransCoder-ST (Rozière et al., 2022) augments TransCoder and DOBF with external unit tests for code translation, while InterCode (Yang et al., 2023d) explores interactive coding with execution feedback. In Section 5.7, we have also shown that the execution results on unit tests serve as natural supervision signals for reinforcement learning on code.
Notably, in March 2023 OpenAI also released an interpreter plugin for ChatGPT999https://openai.com/blog/chatgpt-plugins#code-interpreter, which can accept file inputs from users, generate code according to user instructions, and provide feedback via real-time execution. Zhou et al. (2023a) show that this feature allows GPT-4 to self-debug.
A topic closely related to tool using in LLM researches is planning as intelligent agents, which has been shown to enhance LLMs’ capability both theoretically and empirically (Feng et al., 2023). Ruan et al. (2023) find that LLMs can plan to solve complex tasks using external SQL generators and Python generators, while CodePlan (Bairi et al., 2023) demonstrates they can perform repository-level coding via adaptive planning.
Another stream of works use LLMs to create multi-agent systems for code generation, such as self-collaboration (Dong et al., 2023), ChatDev (Qian et al., 2023), and MetaGPT (Hong et al., 2023). In these frameworks, multiple LLMs are prompted to play distinct roles such as programmer, reviewer, and manager. These roles interact with each other, breakdown code generation into different phases (e.g. designing, coding, testing, and documenting), and collaborate to complete complex tasks.
@@ 7.2 LLMs Integrated into Software Development
Recently, however, the application of code language models have transcended simple code completion. GitHub Copilot is arguably one of the most popular AI code assistants, with diverse features including code generation, vulnerability detection, and license management101010https://github.com/features/copilot, while CodeFuse (Di et al., 2023) also integrates code generation, code translation, code commenting, and testcase generation into a single IDE extension. As code language models become larger, however, their client-side deployment and real-time performance also raise new challenges.
As LLMs continue to advance, building applications on top of them is also evolving into a consequential task itself. Many open-source frameworks for such applications have been released, including LangChain111111https://www.langchain.com/, AutoGPT121212https://github.com/Significant-Gravitas/AutoGPT, and WorkGPT131313https://github.com/team-openpm/workgpt. These frameworks provide abstractions over language models for developers, and are actively revolutionizing the entire process of software development even as this survey is being finalized.
@@ 8 Conclusion and Challenges
- Building code LLM ecosystem for full-life-cycle of software development. While the academia have witnessed an abundance of code models, most have been deployed in the coding stage as IDE plugins while neglecting other stages in the life-cycle of software development. In Section 7.2 we mentioned several inspiring examples, and we are hoping to see more applications of code LMs throughout the full life-cycle of software development, from requirement analysis to DevOps, eventually leading to full-scale ecosystems like those around PyTorch (Paszke et al., 2019) and Hugging Face141414https://huggingface.co/.
@@ Appendix A Benchmarks for Downstrem Tasks
Table 5, 6, 7, 8, 9 list benchmark datasets for code downstream tasks. For the size of these datasets, we use self-reported number whenever available.
@@
Task Date Benchmark Source Size Language Text-to-SQL 1990 ATIS Hemphill et al. (1990); Dahl et al. (1994) 11508 1996 GeoQuery Zelle & Mooney (1996) 877 2000 Restaurants Tang & Mooney (2000) 378 2014-09 MAS Li & Jagadish (2014) 196 2017-02 Yelp Yaghmazadeh et al. (2017) 128 2017-02 IMDb Yaghmazadeh et al. (2017) 131 2017-04 Scholar Iyer et al. (2017) 816 2017-08 WikiSQL Zhong et al. (2017) 80654 2018-06 Advising Finegan-Dollak et al. (2018) 4570 2018-09 Spider Yu et al. (2018c) 10181 2019-06 SParC Yu et al. (2019c) 12726 2019-07 MIMICSQL Wang et al. (2020b) 10000 2019-09 CoSQL Yu et al. (2019b) 15598 2020-05 Criteria-to-SQL Yu et al. (2020) 2003 2020-10 Squall Shi et al. (2020) 11276 2020-10 Spider-Realistic Deng et al. (2021) 508 2021-06 Spider-Syn Gan et al. (2021a) 8034 2021-06 SEDE Hazoom et al. (2021) 12023 2021-06 KaggleDBQA Lee et al. (2021) 400 2021-09 Spider-DK Gan et al. (2021b) 535 2022-05 Spider-SS Gan et al. (2022) 8034 2022-05 Spider-CG Gan et al. (2022) 45599 2023-05 BIRD Li et al. (2023f) 12751 Program Synthesis 2018-02 NL2Bash Lin et al. (2018b) 9305 Bash 2018-08 CONCODE Iyer et al. (2018) 104K Java 2019-10 JuICe Agashe et al. (2019) 1.5M/3725 Python 2021-05 APPS Hendrycks et al. (2021a) 10000 Python 2021-07 HumanEval Chen et al. (2021b) 164 Python 2021-08 MBPP Austin et al. (2021) 974 Python 2021-08 MathQA-Python Austin et al. (2021) 23914 Python 2021-08 PlotCoder Chen et al. (2021c) 40797 Python 2022-01 DSP Chandel et al. (2022) 1119 Python 2022-02 CodeContests Li et al. (2022g) 13610 C++, Python, Java 2022-03 MCoNaLa Wang et al. (2023e) 896 Python 2022-06 AixBench Hao et al. (2022) 336 Java 2022-10 MBXP Athiwaratkun et al. (2023) 12.4K 13 2022-10 Multilingual HumanEval Athiwaratkun et al. (2023) 1.9K 12 2022-10 MathQA-X Athiwaratkun et al. (2023) 5.6K Python, Java, JS 2022-11 ExeDS Huang et al. (2022b) 534 Python 2022-11 DS-1000 Lai et al. (2023) 1000 Python 2022-12 ODEX Wang et al. (2022g) 945 Python 2023-02 CoderEval Yu et al. (2023a) 460 Python, Java 2023-03 xCodeEval Khan et al. (2023) 5.5M 11 2023-03 HumanEval-X Zheng et al. (2023a) 820 Python, C++, Java, JS, Go 2023-06 StudentEval Babe et al. (2023) 1749 Python 2023-06 DotPrompts Agrawal et al. (2023) 10538 Java 2023-09 CodeApex Fu et al. (2023) 476 C++ 2023-09 VerilogEval Liu et al. (2023g) 8645/156 Verilog
@@
Task Date Benchmark Source Size Language Code Translation 2020-06 GeeksforGeeks Rozière et al. (2020) 1.4K C++, Java, Python 2021-02 CodeTrans Lu et al. (2021) 11.8K Java, C# 2021-08 Avatar Ahmad et al. (2023) 9515 Java, Python 2022-06 CoST Zhu et al. (2022b) 132K C++, Java, Python, C#, JS, PHP, C 2022-06 XLCoST Zhu et al. (2022a) 567K C++, Java, Python, C#, JS, PHP, C 2023-03 xCodeEval Khan et al. (2023) 5.6M 11 2023-03 HumanEval-X Zheng et al. (2023a) 1640 Python, C++, Java, JS, Go 2023-08 G-TransEval Jiao et al. (2023) 4000 C++, Java, C#, JS, Python 2023-10 CodeTransOcean Yan et al. (2023) 270.5K 45 Program Repair 2014-07 Defects4J Just et al. (2014) 357 Java 2015-12 ManyBugs Goues et al. (2015) 185 C 2015-12 IntroClass Goues et al. (2015) 998 C 2016-11 BugAID Hanam et al. (2016) 105K JS 2017-02 DeepFix Gupta et al. (2017) 6971 C 2017-05 Codeflaws Tan et al. (2017) 3902 C 2017-10 QuixBugs Lin et al. (2017) 80 Java, Python 2018-05 Bugs.jar Saha et al. (2018) 1158 Java 2018-12 BFP Chakraborty et al. (2022b) 124K Java 2019-01 Bears Madeiral et al. (2019) 251 Java 2019-01 unnamed Tufano et al. (2019a) 21.8K Java 2019-04 BugsJS Gyimesi et al. (2019) 453 JS 2019-05 BugSwarm Tomassi et al. (2019) 1827/1264 Java/Python 2019-05 CPatMiner Nguyen et al. (2019) 17K Java 2019-05 ManySStuBs4J Karampatsis & Sutton (2020) 154K Java 2019-11 Refactory Hu et al. (2019) 1783 Python 2020-07 CoCoNut Lutellier et al. (2020) 24M Java, JS, C, Python 2020-10 Review4Repair Huq et al. (2022) 58021 Java 2020-11 BugsInPy Widyasari et al. (2020) 493 Python 2021-07 TFix Berabi et al. (2021) 105K JS 2021-08 Megadiff Monperrus et al. (2021) 663K Java 2022-01 SSB/TSSB Richter & Wehrheim (2022) 9M/3M Python 2022-10 FixJS Csuvik & Vidács (2022) 324K JS 2022-11 TypeBugs Oh & Oh (2022) 93 Python 2023-03 xCodeEval Khan et al. (2023) 4.7M 11 2023-04 RunBugRun Prenner & Robbes (2023) 450K C, C++, Java, Python, JS, Ruby, Go, PHP 2023-08 HumanEvalPack Muennighoff et al. (2023) 984 Python, JS, Go, Java, C++, Rust
@@
Task Date Benchmark Source Size Language Code Summarization 2016-08 CODE-NN Iyer et al. (2016) 66K/32K C#/SQL 2017-07 unnamed Barone & Sennrich (2017) 150K Python 2018-05 DeepCom Hu et al. (2018a) 588K Java 2018-07 TL-CodeSum Hu et al. (2018b) 411K Java 2018-11 unnamed Wan et al. (2018) 109K Python 2019-02 unnamed LeClair et al. (2019) 2.1M Java 2019-09 CodeSearchNet Husain et al. (2019) 2.3M Go, JS, Python, PHP, Java, Ruby 2023-08 HumanEvalPack Muennighoff et al. (2023) 984 Python, JS, Go, Java, C++, Rust Code Completion 2013-05 GitHub Java Corpus Allamanis & Sutton (2013) 2.1M Java 2016-10 Py150 Raychev et al. (2016a) 150K Python 2016-10 JS150 Raychev et al. (2016a) 150K JS 2023-06 LCC Guo et al. (2023b) 360K Python, Java, C# Commit Message Generation 2017-03 unnamed Jiang & McMillan (2017) 509K Java 2017-04 CommitGen Loyola et al. (2017) 153K Python, C++, Java, JS 2017-08 CommitGen Jiang et al. (2017) 32K/75K Java 2018-09 NNGen Liu et al. (2018) 27K Java 2019-05 PtrGNCMsg Liu et al. (2019c) 64.9K Java 2019-08 CoDiSum Xu et al. (2019a) 90.7K Java 2019-12 ATOM Liu et al. (2022c) 160K Java 2021-05 CommitBERT Jung (2021) 346K Python, PHP, Go, Java, JS, Ruby 2021-07 MCMD Tao et al. (2021) 2.25M Java, C#, C++, Python, JS 2021-07 CoRec Wang et al. (2021b) 107K Java 2023-07 ExGroFi Wang et al. (2023a) 19263 Java 2023-08 CommitChronicle Eliseeva et al. (2023) 10.7M 20 Defect (Vulnerability) Detection 2018-01 CGD Li et al. (2018c) 62K C, C++ 2018-07 Draper VDISC Russell et al. (2018) 12.8M C, C++ 2018-07 SySeVR Li et al. (2018b) 15591 C, C++ 2018-04 unnamed Lin et al. (2018a) 32988 C, C++ 2019-02 unnamed Ponta et al. (2019) 624 Java 2019-09 Devign Zhou et al. (2019) 48687 C 2019-11 unnamed Lin et al. (2021) 170K C, C++ 2019-12 GREAT Hellendoorn et al. (2020) 2.8M Python 2020-01 MVD Zou et al. (2021) 182K C, C++ 2020-02 unnamed Lin et al. (2019) 1471 C 2020-09 ReVeal Chakraborty et al. (2022c) 18K C 2020-09 Big-Vul Fan et al. (2020) 265K C, C++ 2021-02 D2A Zheng et al. (2021) 1.3M C, C++ 2021-05 PyPIBugs Allamanis et al. (2021) 2374 Python 2021-07 CVEfixes Bhandari et al. (2021) 5495 27 2021-08 CrossVul Nikitopoulos et al. (2021) 27476 40+ 2023-04 DiverseVul Chen et al. (2023d) 349K C, C++ 2023-06 VulnPatchPairs Risse & Böhme (2023) 26K C 2023-11 VulBench Gao et al. (2023d) 455 C

### v5

@@ 5.6 Diffusion Models
Currently the Transformer architecture dominates text generation, but several works (Li et al., 2022c; Lin et al., 2023) have also adopted Diffusion Models (Ho et al., 2020) from computer vision for text generation. Recently CodeFusion (Singh et al., 2023) also introduces diffusion models into code modeling, and demonstrates that a 75M diffusion model can outperform StarCoder, CodeT5+, and GPT-3 on three code synthesis datasets.
@@ 5.7 Instruction Finetuning and Reinforcement Learning for Code
In natural language processing, training models on a diverse set of tasks with instruction prefix, known as instruction finetuning, has been shown to unlock the ability of cross-task generalization (Ouyang et al., 2022; Chung et al., 2022; Iyer et al., 2022). At first, these instruction data samples are manually compiled or crowd-sourced (Wei et al., 2022b; Sanh et al., 2022), but later researches find LLM-generated instructions to be sufficient (Wang et al., 2023e; Honovich et al., 2023).
Following these works in natural language, researchers from the code community have applied instruction tuning to their models as well. Wang et al. (2023f) finetune CodeT5+ with 20K instruction data generated by InstructGPT (Ouyang et al., 2022) to obtain InstructCodeT5+. WizardCoder (Luo et al., 2023) follows the methods of WizardLM (Xu et al., 2024) to evolve 20K code Alpaca (Taori et al., 2023) samples into a 78K dataset and uses it to finetune StarCoder. Pangu-Coder 2 (Shen et al., 2023) also uses WizardLM’s Evol-Instruct to generate 68K instruction samples from 20K code Alpaca, but also introduces reinforcement learning via Rank Responses to align Test & Teacher Feedback (RRTF). OctoCoder (Muennighoff et al., 2024), on the other hand, takes a different path and uses Git commit histories as instruction data to finetune StarCoder and CodeGeeX2. More recently, CodeFuse (Liu et al., 2023b) also employs multitask-finetuning and explicitly introduces multiple downstream tasks into their instruction data. The performance of these instruction finetuned code models can also be found in Table 4.
@@
CodeRL (Le et al., 2022) is one such model, which defines four levels of rewards for each generated program (viz. compile error, runtime error, unit test failure, pass) as well as fine-grained token-level reward estimated by a critic model. The actor model, which is an extension of CodeT5, is then trained with REINFORCE algorithm (Williams, 1992). Similarly, CompCoder (Wang et al., 2022c) and PPOCoder (Shojaee et al., 2023) train CodeGPT and CodeT5 respectively with proximal policy optimization (Schulman et al., 2017), while RLTF (Liu et al., 2023d) proposes fine-grained feedback based on the error information and location provided by the compiler, as well as adaptive feedback that takes the ratio of passed test cases into account.
@@ 7.1 LLMs Extended with Coding Tools
Research in the NLP community has shown that LLMs can learn to use external tools such as calculators, MT systems, and search engines (Thoppilan et al., 2022; Schick et al., 2023). As such, interpreter has been used to augment LLMs in complex reasoning tasks. PAL (Gao et al., 2023b) and PoT (Chen et al., 2023c) both extend Codex with Python interpreters for numerical calculations, while ViperGPT (Surís et al., 2023) extends it further by calling vision APIs to extract information from visual input and answer related questions.
However, LLMs do not always produce code that can be interpreted and executed, and there has also been a line of works that explore the possibility of emulating the interpreter with LLMs themselves. Nye et al. (2021) trains LLMs to emulate the execution of programs by outputting program states at each step. More recently, Li et al. (2023a) propose Chain-of-Code, where a real interpreter executes the generated code until an error occurs, whereupon the LLM takes over to simulate execution. Chae et al. (2024) similarly propose Think-and-Execute framework, where an LLM generates pseudo code and then simulates its execution to solve reasoning tasks.
Apart from alleviating the burden of numerical calculation in abstract reasoning tasks, interpreter (together with unit tests) also provides feedback on the process of code generation itself, allowing for interactive generation and refinement of code. Such works include Self-Edit (Zhang et al., 2023c), LeTI (Wang et al., 2023d), OpenCodeInterpreter (Zheng et al., 2024), ProCoder (Bi et al., 2024), and Cycle (Ding et al., 2024), which run model-generated code against unit tests to provide feedback for further refinement. Alternatively, CodeT (Chen et al., 2023a), TiCoder (Bareiß et al., 2022), and CONLINE (He et al., 2024) also utilize the LLM itself to generate unit tests, while Self-Refine (Madaan et al., 2023) sends the generated code to an LLM instead of an interpreter for feedback. Zhou et al. (2023a) show that OpenAI’s interpreter plugin777https://openai.com/blog/chatgpt-plugins#code-interpreter allows GPT-4 to self-debug, while InterCode (Yang et al., 2023d) provides a benchmark for evaluating interactive coding. In Section 5.7 we have also shown that the execution results on unit tests serve as natural supervision signals for reinforcement learning on code.
A topic closely related to tool using in LLM research is planning as intelligent agents, which has been shown to enhance LLMs’ capability both theoretically and empirically (Feng et al., 2023). Ruan et al. (2023) find that LLMs can plan to solve complex tasks using external SQL generators and Python generators, while CodePlan (Bairi et al., 2023) demonstrates they can perform repository-level coding via adaptive planning.
Another stream of works use LLMs to create multi-agent systems for code generation, such as self-collaboration (Dong et al., 2023), ChatDev (Qian et al., 2023), MetaGPT (Hong et al., 2023), LCG (Lin et al., 2024), MAGIS (Tao et al., 2024), and SoA (Ishibashi & Nishimura, 2024). In these frameworks, multiple LLMs are prompted to play distinct roles such as programmer, reviewer, and manager. These roles interact with each other, breakdown code generation into different phases (e.g. designing, coding, testing, and documenting), and collaborate to complete complex tasks.
@@ 7.2 LLMs Integrated into Software Development
Recently, however, the application of code language models have transcended simple code completion. GitHub Copilot is arguably one of the most popular AI code assistants, with diverse features including code generation, vulnerability detection, and license management888https://github.com/features/copilot, while CodeFuse (Di et al., 2023) also integrates code generation, code translation, code commenting, and testcase generation into a single IDE extension. As code language models become larger, however, their client-side deployment and real-time performance also raise new challenges.
As LLMs continue to advance, building applications on top of them is also evolving into a consequential task itself. Many open-source frameworks for such applications have been released, including LangChain999https://www.langchain.com/, AutoGPT101010https://github.com/Significant-Gravitas/AutoGPT, and WorkGPT111111https://github.com/team-openpm/workgpt. These frameworks provide abstractions over language models for developers, and are actively revolutionizing the entire process of software development even as this survey is being finalized.
7.3 Analysis of LLM-Generated Code
As AI code assistants become prevalent, many recent works have also focused on examining AI-generated code from different aspects, including correctness (Nguyen & Nadi, 2022; Yetistiren et al., 2023), bugs (Jesse et al., 2023; Tambon et al., 2024), vulnerabilities (Asare et al., 2023; Sandoval et al., 2023; Perry et al., 2023; Hamer et al., 2024), syntactic robustness (Sarker et al., 2024), efficiency (Niu et al., 2024), and hallucinations (Liu et al., 2024). Asare et al. (2023) and Sandoval et al. (2023)’s studies show that AI code assistants do not introduce extra security risks compared with human programmers, and Hamer et al. (2024) also find that ChatGPT’s responses to security-related questions contain less vulnerabilities than answers from StackOverflow. Contrarily, Perry et al. (2023) find that users write significantly less secure code when assisted by AI. In terms of code complexity, Nguyen & Nadi (2022) find GitHub Copilot to generate low-complexity code with no statistically significant differences between languages, whereas Liu et al. (2023n) find ChatGPT to generate the most complex code in C and the least complex code in Python. Overall, AI code assistants are still in their nascent stage and constantly evolving, and their implications on software development are yet to be investigated systematically.
@@ 8 Conclusion and Challenges
- Building code LLM ecosystem for full-life-cycle of software development. While the academia have witnessed an abundance of code models, most have been deployed in the coding stage as IDE plugins while neglecting other stages in the life-cycle of software development. In Section 7.2 we mentioned several inspiring examples, and we are hoping to see more applications of code LMs throughout the full life-cycle of software development, from requirement analysis to DevOps, eventually leading to full-scale ecosystems like those around PyTorch (Paszke et al., 2019) and Hugging Face121212https://huggingface.co/.
@@ Appendix A Benchmarks for Downstrem Tasks
Table 5, 6, 7, 8, 9, 10 list benchmark datasets for code downstream tasks. For the size of these datasets, we use self-reported number whenever available.
| Task | Date | Benchmark | Source | Size |
| Text-to-SQL | 1990 | ATIS | Hemphill et al. (1990); Dahl et al. (1994) | 11508 |
| 1996 | GeoQuery | Zelle & Mooney (1996) | 877 | |
| 2000 | Restaurants | Tang & Mooney (2000) | 378 | |
| 2014-09 | MAS | Li & Jagadish (2014) | 196 | |
| 2017-02 | Yelp | Yaghmazadeh et al. (2017) | 128 | |
| 2017-02 | IMDb | Yaghmazadeh et al. (2017) | 131 | |
| 2017-04 | Scholar | Iyer et al. (2017) | 816 | |
| 2017-08 | WikiSQL | Zhong et al. (2017) | 80654 | |
| 2018-06 | Advising | Finegan-Dollak et al. (2018) | 4570 | |
| 2018-09 | Spider | Yu et al. (2018c) | 10181 | |
| 2019-06 | SParC | Yu et al. (2019c) | 12726 | |
| 2019-07 | MIMICSQL | Wang et al. (2020b) | 10000 | |
| 2019-09 | CoSQL | Yu et al. (2019b) | 15598 | |
| 2020-05 | Criteria-to-SQL | Yu et al. (2020) | 2003 | |
| 2020-10 | Squall | Shi et al. (2020) | 11276 | |
| 2020-10 | Spider-Realistic | Deng et al. (2021) | 508 | |
| 2021-06 | Spider-Syn | Gan et al. (2021a) | 8034 | |
| 2021-06 | SEDE | Hazoom et al. (2021) | 12023 | |
| 2021-06 | KaggleDBQA | Lee et al. (2021) | 400 | |
| 2021-09 | Spider-DK | Gan et al. (2021b) | 535 | |
| 2022-05 | Spider-SS | Gan et al. (2022) | 8034 | |
| 2022-05 | Spider-CG | Gan et al. (2022) | 45599 | |
| 2023-05 | BIRD | Li et al. (2023g) | 12751 |
@@
Task Date Benchmark Source Size Language Program Synthesis 2018-02 NL2Bash Lin et al. (2018b) 9305 Bash 2018-08 CONCODE Iyer et al. (2018) 104K Java 2019-10 JuICe Agashe et al. (2019) ∗1.5M/3725 Python 2021-05 APPS Hendrycks et al. (2021a) 10000 Python 2021-07 HumanEval Chen et al. (2021b) 164 Python 2021-08 MBPP Austin et al. (2021) 974 Python 2021-08 MathQA-Python Austin et al. (2021) 23914 Python 2021-08 PlotCoder Chen et al. (2021c) 40797 Python 2022-01 DSP Chandel et al. (2022) 1119 Python 2022-02 CodeContests Li et al. (2022g) 13610 C++, Python, Java 2022-03 MCoNaLa Wang et al. (2023g) 896 Python 2022-06 AixBench Hao et al. (2022) 336 Java 2022-10 MBXP Athiwaratkun et al. (2023) 12.4K 13 2022-10 Multilingual HumanEval Athiwaratkun et al. (2023) 1.9K 12 2022-10 MathQA-X Athiwaratkun et al. (2023) 5.6K Python, Java, JS 2022-11 ExeDS Huang et al. (2022b) 534 Python 2022-11 DS-1000 Lai et al. (2023) 1000 Python 2022-12 ODEX Wang et al. (2022g) 945 Python 2023-02 CoderEval Yu et al. (2023a) 460 Python, Java 2023-03 xCodeEval Khan et al. (2023) 5.5M 11 2023-03 HumanEval-X Zheng et al. (2023a) 820 Python, C++, Java, JS, Go 2023-06 StudentEval Babe et al. (2023) †1749 Python 2023-06 DotPrompts Agrawal et al. (2023) ‡10538 Java 2023-09 CodeApex Fu et al. (2023) 476 C++ 2023-09 VerilogEval Liu et al. (2023g) ⋄8645/156 Verilog 2023-11 ML-Bench Liu et al. (2023m) 10040 Bash 2024-01 CRUXEval Gu et al. (2024) 800 Python
@@
Task Date Benchmark Source Size Language Code Translation 2020-06 GeeksforGeeks Rozière et al. (2020) 1.4K C++, Java, Python 2021-02 CodeTrans Lu et al. (2021) 11.8K Java, C# 2021-08 Avatar Ahmad et al. (2023) 9515 Java, Python 2022-06 CoST Zhu et al. (2022b) ∗132K C++, Java, Python, C#, JS, PHP, C 2022-06 XLCoST Zhu et al. (2022a) ∗567K C++, Java, Python, C#, JS, PHP, C 2023-03 xCodeEval Khan et al. (2023) ∗5.6M 11 2023-03 HumanEval-X Zheng et al. (2023a) ∗1640 Python, C++, Java, JS, Go 2023-08 G-TransEval Jiao et al. (2023) ∗4000 C++, Java, C#, JS, Python 2023-10 CodeTransOcean Yan et al. (2023) 270.5K 45 Program Repair 2014-07 Defects4J Just et al. (2014) 357 Java 2015-12 ManyBugs Goues et al. (2015) 185 C 2015-12 IntroClass Goues et al. (2015) 998 C 2016-11 BugAID Hanam et al. (2016) 105K JS 2017-02 DeepFix Gupta et al. (2017) 6971 C 2017-05 Codeflaws Tan et al. (2017) 3902 C 2017-10 QuixBugs Lin et al. (2017) 80 Java, Python 2018-05 Bugs.jar Saha et al. (2018) 1158 Java 2018-12 BFP Chakraborty et al. (2022b) 124K Java 2019-01 Bears Madeiral et al. (2019) 251 Java 2019-01 unnamed Tufano et al. (2019a) †21.8K Java 2019-04 BugsJS Gyimesi et al. (2019) 453 JS 2019-05 BugSwarm Tomassi et al. (2019) 1827/1264 Java/Python 2019-05 CPatMiner Nguyen et al. (2019) †17K Java 2019-05 ManySStuBs4J Karampatsis & Sutton (2020) 154K Java 2019-11 Refactory Hu et al. (2019) 1783 Python 2020-07 CoCoNut Lutellier et al. (2020) 24M Java, JS, C, Python 2020-10 Review4Repair Huq et al. (2022) 58021 Java 2020-11 BugsInPy Widyasari et al. (2020) 493 Python 2021-07 TFix Berabi et al. (2021) 105K JS 2021-08 Megadiff Monperrus et al. (2021) †663K Java 2022-01 SSB/TSSB Richter & Wehrheim (2022) 9M/3M Python 2022-10 FixJS Csuvik & Vidács (2022) 324K JS 2022-11 TypeBugs Oh & Oh (2022) 93 Python 2023-03 xCodeEval Khan et al. (2023) 4.7M 11 2023-04 RunBugRun Prenner & Robbes (2023) 450K C, C++, Java, Python, JS, Ruby, Go, PHP 2023-08 HumanEvalPack Muennighoff et al. (2024) 984 Python, JS, Go, Java, C++, Rust 2024-01 DebugBench Tian et al. (2024) 4253 C++, Java, Python
@@
Task Date Benchmark Source Size Language Code Summarization 2016-08 CODE-NN Iyer et al. (2016) 66K/32K C#/SQL 2017-07 unnamed Barone & Sennrich (2017) 150K Python 2018-05 DeepCom Hu et al. (2018a) 588K Java 2018-07 TL-CodeSum Hu et al. (2018b) 411K Java 2018-11 unnamed Wan et al. (2018) 109K Python 2019-02 unnamed LeClair et al. (2019) 2.1M Java 2019-09 CodeSearchNet Husain et al. (2019) 2.3M Go, JS, Python, PHP, Java, Ruby 2023-08 HumanEvalPack Muennighoff et al. (2024) 984 Python, JS, Go, Java, C++, Rust ∗Code Completion 2013-05 GitHub Java Corpus Allamanis & Sutton (2013) 2.1M Java 2016-10 Py150 Raychev et al. (2016a) 150K Python 2016-10 JS150 Raychev et al. (2016a) 150K JS 2023-06 LCC Guo et al. (2023b) 360K Python, Java, C# Commit Message Generation 2017-03 unnamed Jiang & McMillan (2017) 509K Java 2017-04 CommitGen Loyola et al. (2017) 153K Python, C++, Java, JS 2017-08 CommitGen Jiang et al. (2017) †32K/75K Java 2018-09 NNGen Liu et al. (2018) 27K Java 2019-05 PtrGNCMsg Liu et al. (2019c) 64.9K Java 2019-08 CoDiSum Xu et al. (2019a) 90.7K Java 2019-12 ATOM Liu et al. (2022c) 160K Java 2021-05 CommitBERT Jung (2021) 346K Python, PHP, Go, Java, JS, Ruby 2021-07 MCMD Tao et al. (2021) 2.25M Java, C#, C++, Python, JS 2021-07 CoRec Wang et al. (2021b) 107K Java 2023-07 ExGroFi Wang et al. (2023b) 19263 Java 2023-08 CommitChronicle Eliseeva et al. (2023) 10.7M 20 Defect (Vulnerability) Detection 2018-01 CGD Li et al. (2018b) 62K C, C++ 2018-07 Draper VDISC Russell et al. (2018) 12.8M C, C++ 2018-07 SySeVR Li et al. (2022i) 15591 C, C++ 2018-04 unnamed Lin et al. (2018a) 32988 C, C++ 2019-02 unnamed Ponta et al. (2019) 624 Java 2019-09 Devign Zhou et al. (2019) 48687 C 2019-11 unnamed Lin et al. (2021) 170K C, C++ 2019-12 GREAT Hellendoorn et al. (2020) 2.8M Python 2020-01 MVD Zou et al. (2021) 182K C, C++ 2020-02 unnamed Lin et al. (2019) 1471 C 2020-09 ReVeal Chakraborty et al. (2022c) 18K C 2020-09 Big-Vul Fan et al. (2020) 265K C, C++ 2021-02 D2A Zheng et al. (2021) 1.3M C, C++ 2021-05 PyPIBugs Allamanis et al. (2021) 2374 Python 2021-07 CVEfixes Bhandari et al. (2021) 5495 27 2021-08 CrossVul Nikitopoulos et al. (2021) 27476 40+ 2023-04 DiverseVul Chen et al. (2023e) 349K C, C++ 2023-06 VulnPatchPairs Risse & Böhme (2023) 26K C 2023-11 VulBench Gao et al. (2023d) 455 C